OpenStudy (anonymous):

Are these derivatives correct? f(x)=ln(x^3+3) f' (x)=-2x^2 f(x)=exp(2x^2-1) f^ (x)=exp(2x^2-1)

OpenStudy (bibby):

for the first one, use the chain rule

OpenStudy (anonymous):

Are you sure, isn't the y' of lnx = 1/x?

OpenStudy (bibby):

the derivative of ln(u) = 1/u*u'

OpenStudy (nincompoop):

technically it is 1/x * x but the x is not emphasized as much for some odd reason and that reason is not clear until you learn implicit differentiation for now, your answer couldn't be more wrong

OpenStudy (bibby):

1/x * x ?

OpenStudy (nincompoop):

1/x * x'

OpenStudy (bibby):

o

OpenStudy (nincompoop):

did I miss that?

OpenStudy (bibby):

yeah you didn't have the little apostrophe in the original

OpenStudy (nincompoop):

O I did

OpenStudy (nincompoop):

you're right

OpenStudy (nincompoop):

oh well

OpenStudy (nincompoop):

fkin ey

OpenStudy (nincompoop):

u = x^3 + 3

OpenStudy (bibby):

did you have a drink too much

OpenStudy (nincompoop):

haha just half a bottle of rum